OpenStudy (mathmath333):

At what time between 8-9 will the first right angle will be made by \(\cancel{second}\) minutehand and the hour hand .

OpenStudy (bloomlocke367):

Do you have any ideas?

OpenStudy (anonymous):

Would it be 9 oclock

OpenStudy (mathmate):

Hint: For every 360 degrees the minutes hand moves, the hours hand moves by 30 degrees, so the two hands approaches or diverges by (360-30)/60 = 11/2 degrees per minute.

OpenStudy (mathmath333):

connfused

OpenStudy (mathmate):

 This is the first right angle between the hands between 8 and 9, right?

OpenStudy (mathmath333):

yes

OpenStudy (mathmate):

If the hour hand does not move, we would have a 90 degree at 8:25,agree?

OpenStudy (mathmath333):

yes

OpenStudy (mathmate):

Since the hours hand moves at the rate of 1/12 of the minutes hand, it will take the minutes hand 12/11 of the time (25 minutes) to catch up to the same position of 90 degrees, giving the exact time as 25*12/11=300/11=27 3/11 minutes.

OpenStudy (mathmath333):

yes its correct thnx

ganeshie8 (ganeshie8):

first right angle will be made by `second hand` and the hour hand . is that a typo ? shouldnt it be `minute hand` ?

OpenStudy (mathmath333):

yes lol right

OpenStudy (mathmate):

Right! An oversight! :) Thank you! @ganeshie8

OpenStudy (mathmath333):

@jhonyy9